OpenStudy (anonymous):

Susannah purchased ski equipment for $978 using a six-month deferred payment plan. The interest rate after the introductory period is 21.45%. A down payment of $150 is required as well as a minimum monthly payment of $75. What is the balance at the beginning of the seventh month if only the minimum payment is made during the introductory period?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

can you solve it?

OpenStudy (anonymous):

I think so.... So the bank will flutter you for 21.45 percent of $978 978*0.2145+978=$1187.781 You're paying extra $209.781 right there <--- vacation money right there my friend. anyways 'introductory period' is six months $75 a month 75*6=$450 Subtract that from the adjusted total 1187.781-450=$737.781 is left over to repay the bank after six months. Hope you really had to ski cause I would just save a hundred for 9 pay periods which would be like 4 and a half months. It would take less time to pay for the ski's and you would pay less too. Don't pay the banks.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

I just did...